Cash Budget

Aragon and Associates has found from past experience that 25% of its services are for cash. The remaining 75% are on credit. An aging schedule for accounts receivable reveals the following pattern:

  1. Ten percent of fees on credit are collected in the month that service is rendered.
  2. Sixty percent of fees on credit are collected in the month following service.
  3. Twenty-six percent of fees on credit are collected in the second month following service.
  4. Four percent of fees on credit are never collected.

Fees (on credit) that have not been paid until the second month following performance of the legal service are considered overdue and are subject to a 3% late charge.

Aragon has developed the following forecast of fees:

May $180,000
June 200,000
July 190,000
August 194,000
September 240,000

Required:

Prepare a schedule of cash receipts for August and September. If an amount box does not require an entry, leave it blank or enter "0". Round answers to the nearest dollar.

Aragon and Associates
Schedule of Cash Receipts
For August and September
August September
Cash fees $ $
Received from sales in:
June
July
August
September
Total $ $
Answer #1

Aragon and Associates Schedule of Cash Receipts August September Cash Fees $ 48,500 $ 60,000 Received from Sales in: May June

Working Notes :

Aragon and Associates Schedule of Cash Receipts August =194000*25% September =240000* 25% Cash Fees Received from Sales in: M

Note : As the 4% of the fees is never collected, 3% charge on such is also not considered in calculation, however if we consider the charge then solution would be:

Aragon and Associates Schedule of Cash Receipts August September Cash Fees $ 48,500 $ 60,000 Received from Sales in: May $ 16

In case of any doubts or Issues, please do comment below